6 You are looking at the following information: 5,500 5.5 percent coupon bonds outstanding, $1,000 par Debt: value, 17 yearsMultiple Choice 8.76% polnts 8.22% 8.09% 8.49% 7.99%
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=65*110000
=7150000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*5500*1.04
=5720000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=105*17500
=1837500
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=7150000+5720000+1837500
=14707500
Weight of equity = MV of Equity/MV of firm
Weight of equity = 7150000/14707500
W(E)=0.4861
Weight of debt = MV of Bond/MV of firm
Weight of debt = 5720000/14707500
W(D)=0.3889
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 1837500/14707500
W(PE)=0.1249
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 5 + 1.18 * (6.5)
Cost of equity% = 12.67
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =17x2
1040 =∑ [(5.5*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^17x2
                   k=1
YTM = 5.1441839068
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 5.1441839068*(1-0.33)
= 3.446603217556
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 5/(105)*100
=4.76
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=3.45*0.3889+12.67*0.4861+4.76*0.1249
WACC =8.09%